# TS-SNN: Temporal Shift Module for Spiking Neural Networks

### v1

@@ 2.1 Training of Deep SNNs
Directly learning SNNs is challenging due to the undifferentiable nature of spike firing. The surrogate gradient (SG) approach addresses this issue, thereby enhancing the practicality and applicability of direct SNN training (Neftci et al., 2019; Lee et al., 2020; Wu et al., 2019; Fang et al., 2021a). Recent advancements in SNN architectures have introduced several improvement techniques. MPBN approach (Guo et al., 2023b) incorporates an additional Batch Normalization layer after membrane potential updates to stabilize training processes, DA-LIF(Zhang et al., 2025) introduce an additional learnable decay on conventional LIF. (Duan et al., 2022) proposed the TEBN method, leveraging distinct weights at each timestep to enhance learning dynamics, RSNN(Xu et al., 2024) enhanced spike-based data processing in SNNs. SEW-ResNet (Fang et al., 2021a) introduced a widely utilized ResNet backbone for SNNs, MS-ResNet (Hu et al., 2024) reorganizes the construction of Vanilla ResNet layers. (Zheng et al., 2023) explored spike-based motion estimation for object tracking through bio-inspired unsupervised learning. (Doutsi et al., 2021) introduced a dynamic image quantization mechanism enhancing visual perception quality over time by leveraging both time-SIM and rate-SIM for encoding spike trains. (Xu et al., 2023a, b; Yu et al., 2025) proposed knowledge distillation methods to train deep SNNs using ANNs as the teacher model and SNNs as the student model. These innovations collectively contribute to the advancement of SNNs, making them more practical and effective for a wider range of applications. While existing research addresses various efficient method to enhance SNNs, the survey indicates the absence of an approach fuse features across different timesteps in SNNs.
@@ 2.2 Temporal Shift Modeling
In deep learning, temporal modeling refers to the process of analyzing based on sequential data that evolves over time where the order and timing of data points are important. Due to the intrinsic temporal characteristics of SNNs, there is a natural overlap with video signal processing, suggesting that temporal modeling techniques used in other domains, such as video recognition, can be effectively applied to SNNs. Among traditional temporal modeling approaches, 3D CNNs are the most straightforward method(Guo et al., 2020, 2024). While they can directly process spatiotemporal features, but they come with a high computational cost. To address this, (Lin et al., 2019) proposed TSM, which transforms 3D CNNs into 2D CNNs by shifting a small portion of channels along the temporal dimension, thereby achieving the performance of 3D CNNs at the cost of 2D CNNs. Inspired by the principles behind TSM, ACTION-Net (Wang et al., 2021) introduced a three-path architecture for enhanced action recognition, while the Learnable Gated Temporal Shift Module (LGTSM) (Chang et al., 2019) incorporated learnable kernels to improve temporal fusion efficiency. (Wu et al., 2018a) combined the shift operation with $1\times 1$ convolutions as an efficient alternative to $3\times 3$ convolutions. Further advancements include the proposal of learnable active shifts by researchers (Chen et al., 2019; Jeon & Kim, 2018), and (Liu et al., 2021) implemented a self-attention mechanism through shifting windows to enhance the performance of vision transformers. (Yu et al., 2022) combined spatial shifts with multilayer perceptrons to achieve competitive performance in high-level vision tasks. Additionally, (Li et al., 2023) extended the shift operation to grouped spatial-temporal shifts, demonstrating its applicability to video restoration tasks. These advancements in related studies demonstrate the necessity and effectiveness of temporal modeling. However, current research on SNNs on this domain is still quite limited. Our proposed method aims to fill this gap.
@@ 4.2 Ablation Studies
MPBN(2023b) ResNet-19 1 96.06% ResNet-19 1 78.71%
@@
MPBN(2023b) ResNet-19 10 74.40%
MPBN(2023b) ResNet-20 10 78.70%
@@ A.2 Datasets Details and Augmentations
CIFAR10-DVS (Li et al., 2017) is an event-based version of the CIFAR-10 dataset, captured using a dynamic vision sensor (DVS) camera. This dataset consists of 10,000 event streams, each sized at $128\times 128$, derived from the original CIFAR-10 images. The 10 classes in CIFAR10-DVS each contain 1,000 samples, and the dataset is split into training and testing sets with a 9:1 ratio. During preprocessing, random horizontal flips (with a probability of 0.5) are applied, followed by random selection from augmentations such as rolling, rotation, cutout, and shear. These augmentations enhance the variability and robustness of the dataset(Guan & Zhao, 2022; Guan et al., 2020).
@@ A.3 Experimental Setups
For the CIFAR10-DVS dataset, the initial learning rate was set to 0.1, with a batch size of 32 and 300 training epochs. The parameter $C_{k}$ was set at 32, and $\alpha$ was initialized at 0.2(Guan & Zhao, 2022).

### v3

@@ 2.1 Training of Deep SNNs
Directly learning SNNs is challenging due to the undifferentiable nature of spike firing. The surrogate gradient (SG) approach addresses this issue, thereby enhancing the practicality and applicability of direct SNN training (Neftci et al., 2019; Lee et al., 2020; Wu et al., 2019; Fang et al., 2021a). Recent advancements in SNN architectures have introduced several improvement techniques. MPBN approach (Guo et al., 2023c) incorporates an additional Batch Normalization layer after membrane potential updates to stabilize training processes, DA-LIF (Zhang et al., 2025a) introduce an additional learnable decay on conventional LIF. (Duan et al., 2022) proposed the TEBN method, leveraging distinct weights at each timestep to enhance learning dynamics, RSNN (Xu et al., 2024) enhanced spike-based data processing in SNNs. SEW-ResNet (Fang et al., 2021a) introduced a widely utilized ResNet backbone for SNNs, MS-ResNet (Hu et al., 2024) reorganizes the construction of Vanilla ResNet layers. (Zheng et al., 2023) explored spike-based motion estimation for object tracking through bio-inspired unsupervised learning. (Doutsi et al., 2021) introduced a dynamic image quantization mechanism enhancing visual perception quality over time by leveraging both time-SIM and rate-SIM for encoding spike trains. (Xu et al., 2023a, b; Yu et al., 2025a) proposed knowledge distillation methods to train deep SNNs using ANNs as the teacher model and SNNs as the student model. These innovations collectively contribute to the advancement of SNNs, making them more practical and effective for a wider range of applications. While existing research addresses various efficient method to enhance SNNs, the survey indicates the absence of an approach fuse features across different timesteps in SNNs.
@@ 2.2 Temporal Shift Modeling
In deep learning, temporal modeling refers to the process of analyzing based on sequential data that evolves over time where the order and timing of data points are important. Due to the intrinsic temporal characteristics of SNNs, there is a natural overlap with video signal processing, suggesting that temporal modeling techniques used in other domains, such as video recognition, can be effectively applied to SNNs. Among traditional temporal modeling approaches, 3D CNNs are the most straightforward method(Guo et al., 2020, 2023a). While they can directly process spatiotemporal features, but they come with a high computational cost. To address this, (Lin et al., 2019) proposed TSM, which transforms 3D CNNs into 2D CNNs by shifting a small portion of channels along the temporal dimension, thereby achieving the performance of 3D CNNs at the cost of 2D CNNs. Inspired by the principles behind TSM, ACTION-Net (Wang et al., 2021) introduced a three-path architecture for enhanced action recognition, while the Learnable Gated Temporal Shift Module (LGTSM) (Chang et al., 2019) incorporated learnable kernels to improve temporal fusion efficiency. (Wu et al., 2018a) combined the shift operation with $1\times 1$ convolutions as an efficient alternative to $3\times 3$ convolutions. Further advancements include the proposal of learnable active shifts by researchers (Chen et al., 2019; Jeon & Kim, 2018), and (Liu et al., 2021; Zhang et al., 2025b) implemented a self-attention mechanism through shifting windows to enhance the performance of vision transformers. (Yu et al., 2022) combined spatial shifts with multilayer perceptrons to achieve competitive performance in high-level vision tasks. (Yu et al., 2025b) proposes a spatiotemporal module leveraging frequency characteristics to enhance network performance. Additionally, (Li et al., 2023) extended the shift operation to grouped spatial-temporal shifts, demonstrating its applicability to video restoration tasks. These advancements in related studies demonstrate the necessity and effectiveness of temporal modeling. However, current research on SNNs on this domain is still quite limited. Our proposed method aims to fill this gap.
@@ 4.2 Ablation Studies
MPBN(2023c) ResNet-19 1 96.06% ResNet-19 1 78.71%
@@
MPBN(2023c) ResNet-19 10 74.40%
MPBN(2023c) ResNet-20 10 78.70%
@@ A.2 Datasets Details and Augmentations
CIFAR10-DVS (Li et al., 2017) is an event-based version of the CIFAR-10 dataset, captured using a dynamic vision sensor (DVS) camera. This dataset consists of 10,000 event streams, each sized at $128\times 128$, derived from the original CIFAR-10 images. The 10 classes in CIFAR10-DVS each contain 1,000 samples, and the dataset is split into training and testing sets with a 9:1 ratio. During preprocessing, random horizontal flips (with a probability of 0.5) are applied, followed by random selection from augmentations such as rolling, rotation, cutout, and shear. These augmentations enhance the variability and robustness of the dataset(guan_relative_2022; guan_minimal_2020).
@@ A.3 Experimental Setups
For the CIFAR10-DVS dataset, the initial learning rate was set to 0.1, with a batch size of 32 and 300 training epochs. The parameter $C_{k}$ was set at 32, and $\alpha$ was initialized at 0.2(guan_relative_2022).